OpenStudy (anonymous):

x^2+10=3x-3 . Use the value of discriminant to determine the number and type of roots

OpenStudy (anonymous):

Do you know what the discriminant is?

OpenStudy (zzr0ck3r):

x^2-3x+13 = 0 discriminant = sqrt(b^2-4ac) = sqrt(9-4*1*13) = sqrt(-43) this is complex with imaginary part so two complex solutions

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

First arrange the equation in itz standard form i.e.\[x^{2}+10-3x+3=0=>x^{2}-3x+13=0\] nw waht is discriminant?The discriminant in a quadratic equation is found by the following formula and the discriminant provides critical information regarding the nature of the roots/solutions of any quadratic equation. discriminant= b² − 4ac here a=1 b=-3 c=13 If \[b^{2} − 4ac > 0\]then the equation has two real roots \[b² − 4ac = 0\] then the equation has 1 real root b² − 4ac < 0 then the equation has two complex roots here \[b^{2}-4ac=-43<0\] this means this equation has two complex roots

OpenStudy (anonymous):

@nayyy

OpenStudy (zzr0ck3r):

yes @nayyy discriminant is -43 not sqrt(-43) I was just trying to show you why the -43 causes imaginary roots